# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v25

@@ 3.18.2 Handwritten Text
The work by Karunanayaka et al. [471] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [472]Nilaweera et al. [473] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [474] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [475]Silva et al. [476] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [477]. The study by Walawage and Ranathunga [478] and its follow up study by Walawage [479] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [480] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [481] and Wasalthilake and Kartheeswaran [482], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [483]Ifhaam and Jayalal [484] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [485]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [486] frames their work as an E-Learning platform for hearing-impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognise letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognise Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [403], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [487] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia. The study by eka [488] has used CNN to recognize Sinhala handwritten text. Karunarathne et al. [489] used Gabor-initialised CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.
@@ 3.18.3 Ancient Text
Summarizing optically recognized old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al. [490]. The work of Peiris [491] also focused on OCR for ancient Sinhala inscriptions. Building upon the architecture proposed by Ruwanmini et al. [492], a neural network-based method for recognizing ancient Sinhala inscriptions was proposed by Karunarathne et al. [493]. The study by Wickramarathna and Ranathunga [494] created a system to recognize Brahmi characters, correct errors, and generate Sinhala meanings. Heenkenda and Fernando [495] used Inception-v3 [496], VGG-19 [497], and ResNet-50 [498] to classify Sinhala inscriptions to historical time periods. The work by Gunasekara et al. [499] claims to have created a mobile app that is capable of recognizing and translating early Brahmi characters. The work by Surasinghe and Thanikasalam [500] used CNN models to predict the historical period of ancient Sinhala text. In their subsequent work [501], they extended this work by adding attention to the CNN models.
@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [502]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [503]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [504] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [505] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [130] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conducted by Sheshadri et al. [506] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [507] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [508] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala. Yet another meta-analysis by Halpege [509] compared the commercially available English-Sinhala translation systems provided by Google and Bing through a comprehensive error analysis. This study used academic articles as its domain and ultimately concluded that both translation systems need significant improvements before they can be recommended for widespread usage. The survey by Tafa et al. [510] compared the performance of a number of MT models on Low-Resource languages, including Sinhala.
@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [526] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [527]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [528] which they later extended further [529]Epaliyana et al. [530] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [531] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [532] used a model based on fairseq [533] to improve machine translation between English and Sinhala. The second half of the study by Arafath [400] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [534] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [535]. Nguyen et al. [536] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [537]. Attigala and Weerasinghe [538] conducted an analysis on the effectiveness of ChatGPT [539] in translating Sinhala songs to English. Utsa et al. [540] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [98] and reported better results than mBART [541]. The study by Ranathungaa et al. [542] presents an evaluation of using parallel data from auxiliary domains to enhance EnSI NMT by fine-tuning or further pre-training the models. In their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking, Fernando and Ranathunga [292] used SiTa and EnSi NMT on NLLB and CCAligned as one of the evaluations. Thillainathan et al. [543] proposed to use continual pre-training and intermediate task transfer learning for Fine-tuning what they call multilingual sequence-to-sequence large language models (msLLMs). However, in actuality, what their experiments are on is mBART.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP147147147https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [551] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [552] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [553] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [554]. A rule-based approach for the same was proposed by de Silva and Ahangama [555]. The study by Nanayakkara et al. [556] introduced an English-to-Sinhala transliteration system.
Athukorala and Sumanathilaka [557] proposed a fuzzy logic-based Sinhala transliteration system named Swa-Bhasha. Sumanathilaka et al. [558, 559] then proposed a Trie [560] data structure-based algorithm for word suggestion for the Swa-Bhasha system. An extended analysis of the same work was presented in a later work [561]. A further extension, Swa-Bhasha 2.0, by Dharmasiri and Sumanathilaka [562] used NMT to eliminate word selection ambiguity in the transliteration process. Subsequently, Sumanathilaka et al. [563] claim to have released the Swa-Bhasha Dataset.
@@
The work by Kirov et al. [568] used the Dakshina [157] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [349], ByT5 [569], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [570]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub148148148https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [571] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity. De Mel et al. [572] conducted a comparative analysis for Sinhala transliteration between an extension of the rule-based approach proposed by Tennage et al. [573] and the transliteration as a translation model proposed by Deselaers et al. [574]. They reported across-the-board better performance with the latter. All thier code as well as the test-sets are publicly avaialble149149149https://github.com/kasunw22/Sinhala-Transliterator/. Perera et al. [575] attempted the same task with BERT. Their code is also publicly available150150150https://github.com/Sameera2001Perera/Singlish-Transliterator.
In their short survey discussing Indo-Aryan Languages machine translation and transliteration, Perera and Sumanathilaka [547] covers a number of works [521, 573, 552, 553, 555, 557, 556, 559, 562, 563] on Sinhala transliteration which we have already discussed above.
@@ 3.20 Spelling and Grammar
The open-source data-driven approach proposed by Wasala et al. [607, 608] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [609] attempts to obtain synergy between two algorithms for the same purpose. These efforts [607, 609] were then extended by Subhagya et al. [610]. A rule-based Sinhala spell checker named SinSpell based on Hunspell151151151http://hunspell.github.io/ was introduced by Liyanapathirana et al. [611]. They have also made the tool available152152152http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [612] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [127] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [613]. The study by Udagedara et al. [614] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [615] to correct real-word errors in Sinhala text. Praveen et al. [375] claim to have developed a Sinhala Spell checker with the Random Forest algorithm that archives 82% accuracy.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [78]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [625] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA153153153https://rasa.com/ [626] chat bot. A Sinhala chat bot for train information was proposed by Harshani [627]. Similarly, the tool proposed by Chandrasena et al. [628] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [629] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [630] used off the shelf RASA NLU Engine [626] and Microsoft Bot Network [631] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [632]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [633] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [418] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components. The end product of work by Dasanayaka et al. [634] is a Sinhala chatbot that can answer patient questions on orthopantomography radiology reports which were previously generated by the system itself by analysing dental panoramic tomography (DPT) images.

### v26

@@ 3.18.2 Handwritten Text
The work by Karunanayaka et al. [504] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [505]Nilaweera et al. [506] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [507] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [508]Silva et al. [509] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [510]. The study by Walawage and Ranathunga [511] and its follow up study by Walawage [512] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [513] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [514] and Wasalthilake and Kartheeswaran [515], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [516]Ifhaam and Jayalal [517] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [518]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [519] frames their work as an E-Learning platform for hearing-impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognise letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognise Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [425], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [520] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia. The study by eka [521] has used CNN to recognize Sinhala handwritten text. Karunarathne et al. [522] used Gabor-initialised CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images. Even though the work by Ramanayake and Weerakoon [523] applies OCR on Sinhala handwritten text, its ultimate objective is not recognising the text itself, but the gender of the author. For this, they use several classical ML models as well as GoogleNet and ResNet.
@@ 3.18.3 Ancient Text
Summarising optically recognised old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al. [524]. The work of Peiris [525] also focused on OCR for ancient Sinhala inscriptions. Building upon the architecture proposed by Ruwanmini et al. [526], a neural network-based method for recognising ancient Sinhala inscriptions was proposed by Karunarathne et al. [527]. The study by Wickramarathna and Ranathunga [528] created a system to recognise Brahmi characters, correct errors, and generate Sinhala meanings. Heenkenda and Fernando [529] used Inception-v3 [530], VGG-19 [531], and ResNet-50 [532] to classify Sinhala inscriptions to historical time periods. They later extended their work [533] by proposing to use YOLOv5183183183https://github.com/ultralytics/yolov5 to detect ancient Sinhala inscriptions. The work by Gunasekara et al. [534] claims to have created a mobile app that is capable of recognising and translating early Brahmi characters. The work by Surasinghe and Thanikasalam [535] used CNN models to predict the historical period of ancient Sinhala text. In their subsequent work [536], they extended this work by adding attention to the CNN models.
@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [537]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [538]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [539] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [540] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [133] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conducted by Sheshadri et al. [541] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [542] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [543] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala. Yet another meta-analysis by Halpege [544] compared the commercially available English-Sinhala translation systems provided by Google and Bing through a comprehensive error analysis. This study used academic articles as its domain and ultimately concluded that both translation systems need significant improvements before they can be recommended for widespread usage. The survey by Tafa et al. [545] compared the performance of a number of MT models on Low-Resource languages, including Sinhala. De Silva and Hansadi [546] have also conducted another shallow study on rule-based, statistical, neural, and hybrid MT models for Low-Resource languages.
@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [562] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [563]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [564] which they later extended further [565]Epaliyana et al. [566] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [567] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [568] used a model based on fairseq [569] to improve machine translation between English and Sinhala. The second half of the study by Arafath [422] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [570] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [571]. Later, they published a slightly different version of the work at a different venue but under the same title [572]. Nguyen et al. [573] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [574]. Attigala and Weerasinghe [575] conducted an analysis on the effectiveness of ChatGPT [576] in translating Sinhala songs to English. Utsa et al. [577] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [101] and reported better results than mBART [578]. The study by Ranathunga et al. [579] presents an evaluation of using parallel data from auxiliary domains to enhance EnSI NMT by fine-tuning or further pre-training the models. In their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking, Fernando and Ranathunga [305] used SiTa and EnSi NMT on NLLB and CCAligned as one of the evaluations. Thillainathan et al. [580] proposed to use continual pre-training and intermediate task transfer learning for fine-tuning what they call multilingual sequence-to-sequence large language models (msLLMs). However, in actuality, what their experiments are on is mBART. The work by Zebaze et al. [581] introduces a new NMT pardigm that they call CompTra (compositional translation) which is based on LLMs. They report the results for Sinhala using the FLORES-200 [137] data set, where CompTra applied on NLLB [137] gives the best result.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP184184184https://github.com/google-research/xtreme-up data set created by Ruder et al. [147] contains a Sinhala data set for the transliteration task. Perera and Sumanathilaka [589] introduced a data set185185185https://github.com/Sameera2001Perera/Romanized-Sinhala-Transliteration-Disambiguation-Dataset to disambiguate Sinhala transliteration. The data set contains 600 sentences across 22 ambiguous Romanised Sinhala words.
The early work by Goonetilleke et al. [590] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [591] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [592] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [593]. A rule-based approach for the same was proposed by de Silva and Ahangama [594]. The study by Nanayakkara et al. [595] introduced an English-to-Sinhala transliteration system.
Athukorala and Sumanathilaka [596] proposed a fuzzy logic-based Sinhala transliteration system named Swa-Bhasha. Sumanathilaka et al. [597, 598] then proposed a Trie [599] data structure-based algorithm for word suggestion for the Swa-Bhasha system. An extended analysis of the same work was presented in a later work [600]. A further extension, Swa-Bhasha 2.0, by Dharmasiri and Sumanathilaka [601] used NMT to eliminate word selection ambiguity in the transliteration process. Subsequently, Sumanathilaka et al. [602] claim to have released the Swa-Bhasha Dataset. Swa-bhasha Resource Hub [603] was then created by putting the previous work together for easy access.
@@
The work by Kirov et al. [608] used the Dakshina [160] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [366], ByT5 [609], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [610]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub186186186https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [611] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity. De Mel et al. [612] conducted a comparative analysis for Sinhala transliteration between an extension of the rule-based approach proposed by Tennage et al. [613] and the transliteration as a translation model proposed by Deselaers et al. [614]. They reported across-the-board better performance with the latter. All thier code as well as the test-sets are publicly avaialble187187187https://github.com/kasunw22/Sinhala-Transliterator/. Perera et al. [615] attempted the same task with BERT. Their code is also publicly available188188188https://github.com/Sameera2001Perera/Singlish-Transliterator.
In their short survey discussing Indo-Aryan Languages machine translation and transliteration, Perera and Sumanathilaka [585] covers a number of works [557, 613, 591, 592, 594, 596, 595, 598, 601, 602] on Sinhala transliteration which we have already discussed above. The survey by Deshpande and Deshpande [616] covers the works by Perera et al. [615] and De Mel et al. [612] for Sinhala transliteration, which we have discussed above.
@@ 3.20 Spelling and Grammar
The open-source data-driven approach proposed by Wasala et al. [648, 649] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [650] attempts to obtain synergy between two algorithms for the same purpose. These efforts [648, 650] were then extended by Subhagya et al. [651]. A rule-based Sinhala spell checker named SinSpell based on Hunspell189189189http://hunspell.github.io/ was introduced by Liyanapathirana et al. [652]. They have also made the tool available190190190http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [653] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [130] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [654]. The study by Udagedara et al. [655] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [656] to correct real-word errors in Sinhala text. Praveen et al. [393] claim to have developed a Sinhala Spell checker with the Random Forest algorithm that archives 82% accuracy. [657] introduced LMSpell191191191https://anonymous.4open.science/r/lm-spell-A2D1, an LLM-based toolkit for Sinhala spelling correction.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [81]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [667] where they used, the fasttext model trained by Facebook [123, 124, 125], on a RASA192192192https://rasa.com/ [668] chat bot. A Sinhala chat bot for train information was proposed by Harshani [669]. Similarly, the tool proposed by Chandrasena et al. [670] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [671] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [672] used off the shelf RASA NLU Engine [668] and Microsoft Bot Network [673] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [674]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [675] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [441] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components. The end product of work by Dasanayaka et al. [676] is a Sinhala chatbot that can answer patient questions on orthopantomography radiology reports which were previously generated by the system itself by analysing dental panoramic tomography (DPT) images. Kandamby et al. [677] introduced Cloud Era, a multilingual chatbot utilising LightRAG [678] for agentic behaviour in multi-cloud setups. They claim to have achieved 81% recall for Sinhala. The work by Weerakoon et al. [444] set out to build a Sinhala speech-to-speech chatbot. They select Wav2Vec2-BERT [445] for ASR and utilise a RAG approach for the chatbot. The TTS component is handled with the VITS [397]193193193https://github.com/jaywalnut310/vits model.